Please help me find the sum of this finite geometric series.
2+10+50+250+1250. So far I have a=2, r=5 and n=126. Using the formula
Ssubn=(a(1-4^n))/1-r, I have Ssub126=(2(1-5^126))/1-5. This is where things get out of hand. On my calculator I come up with 5.877x10^87. This is supposed to add up to 1562.

Answer by ptaylor(2198)

YOU ARE RIGHT ON TARGET WITH a and r, BUT:
Where did you get n=126??
The series breaks into: 2x5^0+2x5^1+2x5^2+2x5^3+2x5^4
Try to run the numbers with n=5
Sn=a((r^n)-1)/(r-1)
S=2((5^5)-1)/(5-1)
S=2(3124)/4
S=3124/2=1562
Hope this helps----ptaylor